OpenStudy (anonymous):

The tax on a luxury item in Kansas is computed based upon 2% of the first $1000, 4% of the next $2000 and 5% thereafter. What would the tax be on a luxury item valued at $30,000?

OpenStudy (anonymous):

(2/100*1000)+(4/100*2000)+(5/100*27000) =20+80+1350 =1450